OpenStudy (anonymous):

I'm still having trouble understanding this optimization question: "Find two positive real numbers whose sum is 100 and whose product is a maximum." Help?

OpenStudy (jkristia):

you know you need to find 2 values such that a + b = 100 and a * b gives the max value. You know how to find the max point of a parabola (quadratic equation), so by substituting b with b = 100 = a you can rewrite a * b as a * (100 - a). so now you have f(a) = a(100-a), meaning you have a root at a = 0 and a = 100 0(100-0) = 0 100(100-100) = 0. For a parabola the vertex is the max (or min) value of the function, so if you have a function with the root at a = {0,100}, then you can find your max value at a = ? Makes sense ?

OpenStudy (zzr0ck3r):

x+y = 100 and f(x)=xy but y = 100-x so maximize f(x) = x(100-x) = -x^2+100x take derivative -2x+100 set equal to 0 -2x+100 = 0 x = 50

OpenStudy (zzr0ck3r):

we know this is a max because we have degree 2 and negative leading coefficient.

OpenStudy (zzr0ck3r):

all min/max problems can be written in the form maximize f(x) subject to some constraint. your constraint was given as x+y=100 and you are trying to maximize a product function. But since it has two variables xy we need to write y in terms of x, so we use the constraint. This is like having two equations with two variables and using substitution to solve

OpenStudy (jkristia):

I don't think this is a calculus question, so you cannot take the derivative.

OpenStudy (zzr0ck3r):

why do you say that?

OpenStudy (zzr0ck3r):

well if not, instead of finding the derivative of -x^2+100, use the vertex -b/2a to get the same point, its then the same thing

OpenStudy (zzr0ck3r):

no calc needed and the set up is the exact same thing

OpenStudy (anonymous):

It is a calc problem. It's optimization

OpenStudy (zzr0ck3r):

x+y = 100 and f(x)=xy but y = 100-x so maximize f(x) = x(100-x) = -x^2+100x find vertex -b/2a = -100/2(-1) x = 50

OpenStudy (jkristia):

I just assumed since the question is not marked as a calculus question you should not use the derivative to find the max value.

OpenStudy (zzr0ck3r):

either way, same thing...

OpenStudy (jkristia):

ok - sorry

OpenStudy (zzr0ck3r):

vertex vs derivative are the same at this point

OpenStudy (jkristia):

yes - the problem here is to 'see' how to get the quadratic equation from the given information

OpenStudy (anonymous):

thank you very much!

OpenStudy (zzr0ck3r):

you still need to find y, but use your constraint and this is pretty trivial.

OpenStudy (anonymous):

i've got it... as in I understand now. Thank you again

OpenStudy (zzr0ck3r):

good deal, np

OpenStudy (zzr0ck3r):

when doing other optimization problems, think of this example. It is a great basic concept, and the harder problems are the same, just a little harder to figure out what the constraint is, and what the f(x) is we want to max/min.